# Structure-aware Domain Knowledge Injection for Large Language Models

### v2

@@ 4.2 In-depth Evaluation for Multi-choice QA Application
Table 3: Structure-aware knowledge injection to Llama2 (Touvron et al., 2023b) model series.
@@
Ablation Studies. To further demonstrate the efficacy of each component of our StructTuning method, we conduct a comprehensive ablation study with the English split of the MMedBench dataset. Specifically, we take Llama2-7B as the backbone model, select the English textbooks (Jin et al., 2020) (with 26M tokens) for vanilla and our structure-aware continual pre-training, and use different QA samples for supervised fine-tuning. In Tab. 4, “SFT” refers to vanilla SFT with 10K English QA samples provided in MMedBench’s (Qiu et al., 2024) training split, “SSFT” indicates structure-aware SFT on 10K training data, where the questions are the same as “SFT” while the answers are enhanced with knowledge structure explanation by Llama3-70B, as described in Sec. 3.3. “SSFT*” includes another 8K structure-aware QA syntheses by Llama3-70B, consisting of 18K entries for training. The training hyper-parameters follow the main experiment.
Adaptation English Chinese Japanese French Russian Spanish Average
- SFT 44.54 32.81 26.63 15.27 53.91 42.30 35.91
CPT SFT 46.27 32.57 26.13 17.36 50.00 40.63 35.49
SCPT SFT 46.50 32.14 20.10 18.17 53.91 39.97 35.13
SCPT SSFT 49.96 32.63 22.11 17.52 51.17 41.28 35.78
SCPT SSFT* 49.10 33.92 18.33 27.14 57.42 43.73 38.27
RAG 38.12 29.22 22.61 23.34 53.91 36.47 33.95
In Tab. 4, vanilla CPT+SFT injection paradigm does bring an improvement of 1.73% in the English split, while our SCPT technique with vanilla SFT strategy presents a slightly higher accuracy of 46.50%. However, when combining SSFT with SCPT, our method immediately leads to a significant boost in the English split (49.96% v.s. 44.54%), demonstrating the efficacy and necessity of eliciting the learned structured knowledge to solve practical problems. Moreover, the supplementation of 8K extra QA syntheses (“SSFT*”) surprisingly enhances the model performance on the other five subsets, resulting in a valuable gain on the average accuracy to 38.27%. This result demonstrates the knowledge transferability across different languages (Lai et al., 2023; Qin et al., 2024). After training with SSFT, LLMs can actively utilize the knowledge injected in one language to solve the problem in another language, evidencing our superiority against the traditional SFT technique. In Sec. B.4, we follow Liu et al. (2024b) to randomly generate another 8K SFT pairs for further comparison, where our structure-aware SFT syntheses are verified to better enhance LLMs’ knowledge application.
In addition, we observe that the commonly used RAG (Lewis et al., 2020) strategy does not bring significant advantages to the MMedBench evaluation. The main reason lies in the gap between the pre-training corpus (comprising official knowledge statements from textbooks) and evaluated QA samples (originating from practical diagnosis records). Knowledge injection by (S)CPT and (S)SFT shows more advantages in this situation.
5 Conclusion
This work pioneers in incorporating structure-aware methodologies to enhance domain knowledge injection into large language models. Through a novel SCPT-SSFT paradigm, we have set a new precedent for adapting LLMs to specialized domains, and the promising and scalable results underscore the viability and potential of our method. We hope to inspire further research in efficient and effective domain adaptation, moving a step closer to models that can truly emulate human intelligence.
Limitation. A noteworthy limitation of our work is the under-explored full potential of our StructTuning strategy for knowledge injection. Although we have estimated a scaling law in Eq. 3, the data requirement to achieve 100% of the desired effectiveness remains unverified. Besides, our method introduces negligible data preprocessing costs to extract the domain knowledge structure. However, Sec. B.2 indicates our method can still reduce the overall adaptation cost to 10% since only a small part of training corpus is required. We will delve into the investigations in our future work.
Appendix A Implementation Details
A.1 Detailed Setup on LongBench
To focus on the investigation of knowledge injection, we choose 7 subsets from LongBench (Bai et al., 2023) across single- and multi-document QA tasks in English and Chinese, and the remaining synthetic or code-orientated tasks are eliminated:
  • Single-Doc QA. For single-document QA, we take three subsets from LongBench: (1) Qasper (Dasigi et al., 2021), featured by question-answering over NLP technical papers and annotated by NLP practitioners; (2) MultiFieldQA (Bai et al., 2023), manually curated from multiple data sources and annotated by Ph.D. students; and (3) MultiFieldQA-zh, the Chinese split also provided by Bai et al. (2023), covering multiple Chinese scenarios. MultiFieldQA contains 150 Context-Question-Answer triplets to test, and the others adopted subsets include 200 pieces of test samples respectively.
  • Multi-Doc QA. Multi-document QA requires LLMs to extract and combine information from multiple documents to derive the answer, which is generally more challenging than single-doc QA. We take four multi-hop QA datasets: (1) HotpotQA (Yang et al., 2018), containing 2-hop questions written by native speakers given two related paragraphs; (2) 2WikiMultihopQA (Ho et al., 2020), involving up to 5-hop questions synthesized through manually designed templates on Wikipedia passages; (3) MuSiQue (Trivedi et al., 2022), carefully composed with up to 4-hop reasoning on an increased number of supporting and distracting context evidence; and (4) Dureader (He et al., 2017), developed based on Baidu Search and Baidu Zhidao and filtered by Bai et al. (2023) to reduce the data noise. Each subset has 200 test samples.
For data entries in Single-Doc QA, we extract the knowledge structure for each single passage; in Multi-Doc QA, we identify the knowledge structure across multiple passages for each test sample. There are ultimate 1350 question-answer-passage(s)-(knowledge)structure quadruples to evaluate knowledge injection approaches on LongBench.
A.2 Detailed Setup on MMedBench
Table A1: Sample statistics on MMedBench.
Split English Chinese Japanese French Russian Spanish Total
Train 10,178 27,400 1,590 2,171 1,052 2,657 45,048
Test 1,273 3,426 199 622 256 2,742 8,518
Data for Continual Pre-Training. To investigate high-quality domain knowledge injection for LLMs, we collect 18 English textbooks and 33 Chinese textbooks from the National Medical Board Examination in the USA and Mainland China, respectively (Jin et al., 2020). All collected textbooks are originally in PDF format and Jin et al. (2020) converted them into digital text via OCR and performed some clean-up pre-processing strategies to reduce the data noise. The English and Chinese textbooks count for around 26.1M and 21.5M tokens by Llama2 tokenizer (Touvron et al., 2023b). Then, we randomly sample an extra 28M textbook corpora from MMedC (Qiu et al., 2024) for the other languages (except the unavailable Japanese textbooks), and the statistics are displayed in Tab. A2.
Table A2: Sample statistics on training data. $\dagger$” and $\ddagger$” refer to different sizes.
Stage English Chinese Japanese French Russian Spanish Total
CPT 26.1M 21.5M - 8.1M 10.3M 10.1M 76.1M
[…]
SFT 18.8K 39.1K 1.6K 5.3K 5.9K 7.5K 78.2K
[…]
Data for Supervised Fine-Tuning. As introduced in Sec. 3.3, we prompt Llama3-70B (Dubey et al., 2024) to build the structure-aware answer explanations on top of the raw SFT samples in MMedBench’s training split, and generate extra QA pairs by traversing the extracted knowledge structure from textbooks. The final quantity statistics are presented in Tab. A2.
A.3 Resource Requirement
Appendix B Additional Experiments
B.1 Ablation on Knowledge Structure Extraction
Extracting domain knowledge structure is a prerequisite for subsequent knowledge injection (including both SCPT and SSFT) for language models. In Sec. 3.1, we propose a bottom-up strategy to re-chunk the texts from domain textbooks, summarize a title for each chunk, and send the title list to a specialized 7B model to derive the knowledge structure. The prompt template is displayed in Appendix C.
In fact, we argue that due to the language diversity, a perfectly recovered table of contents of textbooks is unnecessary for domain knowledge injection. A reasonable knowledge structure is sufficient enough. In Tab. A3, we individually adopt few-shot GPT-3.5-Turbo (Brown et al., 2020) and LLaMA3-70B (Dubey et al., 2024) models to extract medical knowledge structure from 18 English textbooks (Jin et al., 2020) (with 26.1M tokens) for subsequent knowledge injection (the backbone LLM is LLaMA2-7B (Touvron et al., 2023b)). Although they present 3.6%-3.7% enhancement on MMedBench (Qiu et al., 2024)’s English test set (denoted as “improvement”), leveraging GPT-3.5-Turbo and LLaMA3-70B is either expensive or time-consuming. GPT-3.5-Turbo costs around 15 dollars to process 26M tokens, while LLaMA3-70B takes around 1.5 hours on 2 A100-80G GPUs, of which both limit the scalability of data pre-processing for structure-aware knowledge injection.
Model Improvement Time Consumption Extra Cost
GPT-3.5-Turbo +3.58 - 15\$
LLaMA3-70B +3.72 1.5h -
Ours-7B +3.69 0.2h -
@@ B.1 Ablation on Knowledge Structure Extraction
Inspired by Liu et al. (2024a), we distilled the knowledge structure extraction capability from giant LLMs to a LLaMA2-7B model via supervised fine-tuning. In particular, we instruct LLaMA3-70B to generate 22K training examples (pairs of raw knowledge points and extracted knowledge structures) from Wikipedia, and train a LLaMA2-7B model at a batch size of 128 and a learning rate of 2e-5 for 1 epoch. After utilizing the specialized 7B model to identify the knowledge structure in medical textbooks, as shown in Tab. A3, the results translate to comparable performance on structure-aware knowledge injection. Meanwhile, the inference cost significantly decreases to 0.3 hours, which is more scalable to handle a larger amount of domain corpus.
B.2 Comparison on Training Costs
Table A4: Comparison of training costs for knowledge injection.
Paradigm Corpus Improvement Pre-process Total
@@ B.2 Comparison on Training Costs
B.3 F1-Score Evaluation on LongBench
In Sec. 4.1, we mainly follow Zhu & Li (2023b) to investigate the memorization and understanding of injected knowledge by calculating the knowledge recall in models’ responses. Here we report the F1-score measure over the Open-Book QA (OBQA) and Closed-Book QA (CBQA) settings for a thorough comparison. Note that here we use the vanilla question prompt to obtain concise answers, instead of the CoT prompt used in Sec. 4.1 to elicit models’ memorized knowledge. The evaluated models are the same as Sec. 4.1.
In Tab. A5, we report the Open-Book QA (OBQA) baseline for Llama2-7B with passages as inputs, which shows the best performance on MultiFieldQA (MFQA) (Bai et al., 2023) and 2WikiMultihopQA (2Wiki) (Ho et al., 2020) subsets. Then, we establish the Closed-Book QA (CBQA) baseline by traditional CPT+SFT to inject passage contents into model parameters, and supplement the experiment of our SCPT+SSFT technique for comparison. According to the results shown in Tab. A5, CPT+SFT slightly improves the QA performance on several subsets (such as MultiFieldQA-zh (MFQAzh) (Bai et al., 2023)), while the overall F1-Score measure is still inferior to the OBQA performance. In contrast, our SCPT+SSFT approach successfully boosts the closed-book QA performance to 20.1% on average, even surpassing the open-book QA baseline of 18.7%.
The results are consistent with Sec. 4.1, which jointly demonstrate the effectiveness of structure-aware knowledge injection for large language models.
Table A5: F1 Score evaluation of Open-Book QA (OBQA) and Closed-Book QA (CBQA) tasks on the LongBench (Bai et al., 2023) dataset. The best results are marked in bold, and the secondary results are marked with underlines. The backbone model is Llama2-7B (Touvron et al., 2023b).
Task Adaptation SingleDoc-QA MultiDoc-QA Average
Qasper MFQA MFQAzh HpQA 2Wiki Musiq Duzh
OBQA - 19.2 36.8 11.9 25.4 32.8 9.4 5.2 18.7
CBQA CPT+SFT 16.8 23.1 13.2 21.3 19.1 10.4 13.4 16.8
SCPT+SFT 15.2 21.5 15.2 14.9 19.8 5.6 14.1 15.2
SCPT+SSFT 19.7 23.5 19.5 26.4 24.1 12.2 15.4 20.1
B.4 Ablation on SFT Data Synthesis
@@
Appendix C Prompt Template for Knowledge Structure Extraction

### v3

@@ B.1 Knowledge Structure Extraction
Table A6: Structure-aware knowledge injection to Llama2 (Touvron et al., 2023b) model series.
@@
Inspired by Liu et al. (2024a), we distilled the knowledge structure extraction capability from giant LLMs to a LLaMA2-7B model via SFT. In particular, we instruct LLaMA3-70B to generate 22K training examples (pairs of raw knowledge points and extracted knowledge structures) from Wikipedia, and train a LLaMA2-7B model at a batch size of 128 and a learning rate of 2e-5 for 1 epoch. After utilizing the specialized 7B model to identify the knowledge structure in medical textbooks, as shown in Tab. A5, the results translate to comparable performance on structure-aware knowledge injection. Meanwhile, the inference cost significantly decreases to 0.3 hours, which is more scalable to handle a larger domain corpus.
Table A7: Model evaluation on general benchmarks.
Injection MMLU C-Eval TruthfulQA WinoGrande ARC_c
Before 0.46 0.35 0.49 0.52 0.51
After 0.43 0.35 0.43 0.51 0.56
B.2 Evaluation of Approach’s Generalization
In addition to the Llama3 models in previous experiments, we also investigate the generalization ability of our SCPT+SSFT paradigm on Llama2 (Touvron et al., 2023b) and InternLM2 (Zheng et al., 2024) model series. As shown in Tab. A6, our method leads to consistently significant improvements on InternLM2-7B (+4.46%), Llama2-7B (+8.78%) and Llama2-13B (+6.17%) backbone models. The results further demonstrate the generalizability and scalability of our StructTuning strategy across model architectures and sizes.
B.3 Investigation on Model Overfitting
Here we provide a supplementary evaluation on common benchmarks in Tab. A7, which indicates the current methodology of structure-aware knowledge injection does not significantly hurt LLM’s general capabilities, and even brings slight enhancement to the ARC_c benchmark (maybe the knowledge structure enhances the reasoning ability). As previous research states, the issue of overfitting can be further mitigated by incorporating common QA examples, and we leave this in our future work.
B.4 Comparison on Training Costs
Table A8: Training costs for knowledge injection.
Paradigm Corpus Improve. Preprocess Total
@@ B.4 Comparison on Training Costs
Table A9: Ablation studies of SCPT on MMedBench subsets. The base model is Llama2-7B.
Adaptation English Chinese Japanese French Russian Spanish Average
CPT+SFT 46.27 32.57 26.13 17.36 50.00 40.63 35.49
Ours-FixTmpl1 48.27 32.86 20.61 23.70 56.17 42.56 37.36
Ours-FixTmpl2 48.10 32.99 21.23 23.97 55.97 43.10 37.56
Ours-RemoveL1 47.90 32.90 20.11 24.63 57.10 43.43 37.68
Ours-RemoveL2 48.05 33.63 21.62 24.11 57.49 43.13 38.00
Ours-RemoveL3 48.47 33.14 20.15 23.33 56.86 43.65 37.60
Ours-NTPLoss 48.99 33.15 20.57 25.31 56.78 42.94 37.96
Ours-Full 49.10 33.92 18.33 27.14 57.42 43.73 38.27
B.5 Ablation on Structured Knowledge Injection
During the Structure-aware Continual Pre-Training (SCPT) stage, we proposed to learn specific text chunks (knowledge points) in the condition of the mindmap inputs (knowledge structures), in order to relate the knowledge points to corresponding structure nodes. In this section, we conduct a series of ablation studies to investigate the design efficacy. The vanilla CPT+SFT paradigm is adopted as the comparison baseline, where the Llama2-7B model is trained with CPT and SFT data on the English subset of MMedBench, while tested on all subsets across six languages. The hyper-parameter settings follow the main experiment in our manuscript. The empirical results are presented in Tab. A9.
First, we investigate the choice of formatting template to convert the knowledge structure to a mindmap condition. In particular, we try to fix the template to convert all knowledge mindmaps for SCPT, and randomly select two templates to repeat the experiment. According to Tab. A9, fixed SCPT templates lead to inferior performance against randomly choosing the template from the diversified 20 template pool. This is consistent with Zhu and Li (2023a)’s observation, that text rewriting can provide better knowledge augmentation for large language models.
Then, we explore the impact of the extracted knowledge structure itself. In MMedBench, a 3-layer knowledge structure (follow the chapter-section-subsection hierarchy) is constructed for each textbook, and we respectively remove the 1st (chapter), 2nd (section), and 3rd (subsection) layer of the hierarchy during knowledge injection. As Tab. A9 shows, removing the top layer (chapter) leads to the worst performance of 47.90%, because the remaining knowledge points cannot effectively relate to each other without the organization of the top layer. On the other hand, removing the bottom layer (subsection) performs slightly better on the English subset (because of the controlled structure-information lost), but hinders the cross-language knowledge utilization on the remaining subsets (e.g., 37.60% on average across six languages).
Finally, we revisit the modeling choice of the mindmap-conditioning learning. Specifically, we try to turn the conditional modeling $p({\bm{x}}^{k}|{\bm{s}}^{k})$ back to complete next-token prediction $p({\bm{x}}^{k},{\bm{s}}^{k})$ (the next-token prediction loss is computed on mindmap condition as well). According to Tab. A9, the performance is slightly inferior to our full version of SCPT strategy (e.g., 37.96% v.s. 38.27% on Average). Therefore, we reserve conditional modeling for our SCPT stage.
B.6 Ablation on SFT Data Synthesis
@@
B.7 In-depth Comparison on Retrieval-Augmented Generation
Table A11: Ablation on the hyper-parameter settings for the RAG baseline.
ChunkSize 256 512 1024
RetrieveNum 10 5 3 5 3 2 3 2 1
Accuracy 35.08 37.67 38.04 36.42 38.12 37.99 35.00 36.89 38.07
Table A12: Attempts to integrate hybrid-search and reranker models.
Hybrid Reranker Accuracy
$\times$ $\times$ 38.12
$\surd$ $\times$ 37.97
$\times$ $\surd$ 37.52
$\surd$ $\surd$ 37.75
In Sec. 4.2, we briefly compare RAG adaptation and injection-based approaches in the MMedBench dataset, and this section provides more implementation details and further investigations on the popular retrieval-augmented generation approach.
Experimental Settings. On the implementation of the RAG baseline, we utilize the BAAI/bge-m3 (Chen et al., 2024) embedding model for dense retrieval, due to its state-of-the-art and multi-lingual semantic retrieval ability. For the experiments in Tab. 4, we take the same 26M English CPT data as the knowledge base, re-chunk the data corpus for every 512 tokens, and retrieve top-3 related chunks as context inputs for LLM’s generation process. The retrieval process is implemented using the LlamaIndex [^2] https://www.llamaindex.ai/ framework.
Additional Experiments. We also conduct a variety of experiments to evaluate the hyper-parameters for the RAG baseline. As shown in Tab. A11, changing the chunk size and retrieved chunk number cannot bring any significant benefits. The core reason lies in the gap between user query and retrieved chunks. In particular, user queries contain many descriptive and quantitative sentences and numbers (such as the example in Fig. A3, “They enrolled 800 patients in the study, half of which have breast cancer”.), and may even talk about an entirely new thing that has not been recorded in the knowledge base.
Furthermore, we also try to use the hybrid (dense+sparse) search strategy and larger rerank model (BAAI/bge-reranker-v2-m3 (Chen et al., 2024)) to enhance the retrieval quality. However according to the results in Tab. A12, the semantic gap between user queries and retrieved chunks still exists. Introducing the hybrid search and rerank model even gets worse performance (e.g., the keyword age may be considered a key factor for hybrid search, but it cannot help to derive the answer of test sensitivity).
Conclusion. RAG may assist in some knowledge-intensive tasks for information-seeking, but will encounter problems when there exists a significant semantic gap between user query and retrieved documents. MMedBench is a typical scenario, where LLMs are asked to derive medical diagnoses with proper reasoning according to the descriptions of patients or medical examinations. In this case, the retrieval process introduces too many unrelated chunks and hurts LLM’s QA judgments. Fig. A3 provides an example where the retrieved chunks are actually unrelated to the complicated user query (the user asks about the analysis of a given research study, but the retrieved documents contain several keywords, e.g., age, while having nothing to do with the blood test study.)
Figure A3: An example of retrieved document/chunk based on a given query.
B.8 Knowledge Memorization on LongBench
Here, we also use lexical ROUGE-L (Lin, 2004) and semantic BERTScore (Zhang et al., 2020) to quantify the memorization of injected knowledge structures, by comparing the mindmap in models’ responses (as Fig. 5 displays) with ground-truth answers. The results in Tab. A13 indicate a relatively good memorization of the injected knowledge mindmap, emphasizing the efficacy of our SCPT strategy.
Table A13: MindMap Recall
F1-Score BERTScore
0.61 0.87
B.9 F1-Score Evaluation on LongBench
Table A14: F1 Score evaluation of Closed-Book QA (CBQA) tasks on the LongBench (Bai et al., 2023b) dataset. The best results are marked in bold, and the secondary results are marked with underlines. The backbone model is Llama2-7B (Touvron et al., 2023b).
Task Adaptation SingleDoc-QA MultiDoc-QA Average
Qasper MFQA MFQAzh HpQA 2Wiki Musiq Duzh
CBQA CPT+SFT 16.8 23.1 13.2 21.3 19.1 10.4 13.4 16.8
SCPT+SFT 15.2 21.5 15.2 14.9 19.8 5.6 14.1 15.2
SCPT+SSFT 19.7 23.5 19.5 26.4 24.1 12.2 15.4 20.1
In Sec. 4.1, we mainly follow Zhu and Li (2023b) to investigate the memorization and understanding of injected knowledge by calculating the knowledge recall in models’ responses. Here we report the F1-score measure over the Closed-Book QA (CBQA) settings for a thorough comparison. Note that here we use the vanilla question prompt to obtain concise answers, instead of the CoT prompt used in Sec. 4.1 to elicit models’ memorized knowledge. The evaluated models are the same as Sec. 4.1.
In Tab. A14, we report the Closed-Book QA (CBQA) baseline by traditional CPT+SFT to inject passage contents into model parameters, and supplement the experiment of our SCPT+SSFT technique for comparison. According to the results shown in Tab. A14, our SCPT+SSFT approach successfully boosts the closed-book QA performance to 20.1% on average. The results are consistent with Sec. 4.1, which jointly demonstrate the effectiveness of structure-aware knowledge injection for large language models.
Appendix C Detailed Related Work
Domain Adaptation for LLMs. While pre-trained LLMs possess promising capabilities, their performance is often hampered by the scope and recency of their training data, which particularly affects smaller models in downstream applications (Zhao et al., 2023; Wang et al., 2023a). Continual Pre-Training (CPT) addresses this by perpetually updating a pre-trained model with domain-specific content (Sun et al., 2020; Xu et al., 2023b). , with parameter-efficient tuning methods devised to curtail training costs (Hu et al., 2021; Liu et al., 2024c). To keep pace with the latest information, models can be fine-tuned with supervised instruction-response pairs (SFT), thus staying current with the advancing knowledge landscape (Mecklenburg et al., 2024; Qiu et al., 2024). Existing literature confirms that combining CPT and SFT is effective for LLMs to remain precise and up-to-date in dynamic fields like medicine (Wang et al., 2023b; Qiu et al., 2024) and coding (Roziere et al., 2023; Guo et al., 2024). Our study builds upon this CPT-SFT framework, innovating with SCPT-SSFT strategies to efficiently and effectively infuse domain knowledge with the inherent structure hierarchy.
Conditional Language Modeling. The idea of continual pre-training language models on domain corpus in the condition of the knowledge structure is mainly inspired by CTRL (Keskar et al., 2019). Keskar et al. (2019) demonstrates the effectiveness of steering text generation through control codes (one or two words) that signify the desired genre, style, or task. In the era of LLM, system prompt plays a similar role in controlling models’ responses to adapt to different needs and functionalities, such as role-playing, language style transfer, task setting, and behavior setting (Brown et al., 2020; Wang et al., 2023d; Bai et al., 2023a). Our SCPT approach extends the control codes or system prompts to domain-specific knowledge structures, so as to guide the learning process and tailor the model’s output more closely to specialized fields.
Structure-aware Knowledge Aggregation. Knowledge structure has been widely explored in the recent LLM community. In conventional paradigms, researchers extract entity-relation-entity triplets from texts to construct knowledge graphs (Pan et al., 2024), to enhance LLMs’s factual knowledge and logical reasoning by feature aggregation (Liu et al., 2020; Zhang et al., 2022), prompt engineering (Wen et al., 2023; Wang et al., 2023c), information searching (Logan IV et al., 2019), training data synthesis (Tang et al., 2024), etc. In these cases, each node corresponds to either a specific entity or an abstract concept, lacking the capability to present an informative and self-contained knowledge point. Some works have recently related a piece of descriptive text to a knowledge point, and constructed the knowledge structure for LLMs’ retrieval-augmented generation (Sarthi et al., 2024; Dong et al., 2024), where the top-to-down retrieval provides precise information-seeking paths along the knowledge structure. In this paper, we extend the structure-aware knowledge aggregation to LLMs’ training phase, injecting the whole domain knowledge structure into LLMs’ by linking training samples to corresponding knowledge points and reasoning paths.
Data Augmentation and Synthesis. Due to the lack of high-quality datasets, data augmentation has emerged as a promising solution to mimic real-world patterns (Liu et al., 2024b). Traditional methods aim to artificially expand the training dataset size (Xu et al., 2023a; Mukherjee et al., 2023) or generate entirely new samples that could help models learn better or adapt to specific tasks (Tang et al., 2024). Yet, they often overlook the structured nature of domain knowledge, and the aimlessly generated samples may also lack diversity (Ovadia et al., 2023; Mecklenburg et al., 2024), leading to potentially suboptimal training outcomes when applied for domain adaptations (Mecklenburg et al., 2024; Tang et al., 2024). By contrast, our SSFT design is an innovative departure to address the challenge of retaining and utilizing the structured knowledge inherent in domain-specific content.
Appendix D Prompt Template for Knowledge Structure Extraction